40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Oct 29 15:34:36 2010, Zach & Koji, Laser, GYRO, CCW loop characterization bart-simpson-generator.png
    Reply  Sat Oct 30 20:02:01 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (1) open_loop_primary.pdf
       Reply  Sat Oct 30 20:18:06 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (2) error_primary.pdffeedback_primary.pdf
          Reply  Sat Oct 30 20:57:19 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (3) error_secondary.pdf
             Reply  Sat Oct 30 22:49:04 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (4) 
                Reply  Sat Oct 30 23:41:31 2010, Koji, Laser, GYRO, gyro characterization Oct 27, 2010 (5) 
Message ID: 1119     Entry time: Sat Oct 30 20:57:19 2010     In reply to: 1118     Reply to this: 1120
Author: Koji 
Type: Laser 
Category: GYRO 
Subject: gyro characterization Oct 27, 2010 (3) 

[Zach / Koji]

Measurement of the secondary (CW) cavity error signal.

Method:

- Lock the primary cavity.
- Tune the VCO frequency such that we can transmit the secondary (CW) beam to the transmission CCD. Tune the error signal to zero.
- Leave the secondary loop open. Look at the error signal of it in order to determine whether it is in the linear range or not. (Yes it was)
- Measure the error signal while the loop is open or closed.

Result:

- The control bandwidth seems to be ~100Hz. The control gain seems to be ~3.
- The error signal seems to be completely dominated by the dark noise of the detection system (= PD+demodulator).

Thought:

- The gain and the bandwidth are way too low to obtain the gyro signal from the VCO feedback
- If you look at the structures above 100Hz, the optical gain is about ten times smaller than that of the primary loop.
  What the heck is and the bandwidth are way too low to obtain the gyro signal from the VCO feedback

- Note that the signal is supposed to be amplified by a factor of +41. This means the dark noise floor level is ~25nV/rtHz.

To Do:

  • Calibrate these spectra in Hz/rtHz.
    • How much is the optical gain in V/Hz or V/m.
    • Overlay calibrated spectra of the primary and secondary loops.
    • There looks excess noise below 10Hz. We must reveal what it is through the measurements.
  • This error signal is the gyro output signal as far as there is no feedback.
    Calibrate this signal to (rad/s)/rtHz (or rad/Hz). Put this measurement on the noise budget plot.
  • Convert the dark noise measurement into the noise budget of the Gyro signal.
  • Measure the control loop gain. Investigate the shape of the loop up to 100kHz.
     
  • Measure the shot noise level
    • How much DC power do we typically have? (Actually we don't have the record during this measurement, but we can measure it again)
  • Measure low frequency spectra with CDS
  • How to improve the noise floor
    • Understand why the optical gain is so low.
    • Is this noise level disturbing in terms of the gyro requirement? How about in the low frequency?
    • How much is the demodulator noise? Put a 50-ohm terminator on the cable instead of the PD. Then measure the same signal.
    • How much is the gain of the input stage. It is supposed to be +41, but we are not sure.
      => Think about the noise level at the demodulator output.
    • How much is the noise level of the PD measured by an RF analyzer? Is it consistent with the above analysis?
    • Do we need a new resonant RF PD? How much should we improve the noise level with the new PD? And how much noise does the new one actually have?
  • Where the VCO frequency noise comes? ==> Measurement of the VCO noise.

 

Attachment 1: error_secondary.pdf  92 kB  | Hide | Hide all
error_secondary.pdf
ELOG V3.1.3-